Using Gauss’ Theorem calculate the flux of the vector field F ˆ
i ˆ
j kˆ = x + y + z r through
the surface of a cylinder of radius A and height H, which has its axis along the z-axis
and the base of the cylinder is on the xy-plane.
1
Expert's answer
2019-03-15T12:47:51-0400

Gauss' theorem states

"\rm{Flux}=\int \vec F d\vec A=\intop \rm{div}\vec F dV"

Since

"\rm{div}\vec F=\frac{\partial F_x}{\partial x}+\frac{\partial F_y}{\partial y}+\frac{\partial F_z}{\partial z}=\frac{\partial x}{\partial x}+\frac{\partial y}{\partial y}+\frac{\partial z}{\partial z}=3"

we obtain

"\rm{Flux}=\intop \rm{div}\vec F dV=\intop 3 dV=3V=3 \pi A^2 H"